# Digital FAST: An AI-Driven Multimodal Framework for Rapid and Early Stroke Screening

### v1

@@ V-A Implementation Details
As illustrated in Fig. II, the proposed system consists of multiple modality specific branches designed to process heterogeneous input signals. Facial landmark sequences were modeled using a Transformer encoder initialized from a pretrained Audio Spectrogram Transformer (AST) backbone with a base configuration, while speech signals were processed using an AST model with a tiny configuration to extract discriminative acoustic representations from mel spectrograms. For upper-body pose data, an MLP-Mixer architecture was employed and trained from scratch to capture temporal and spatial interactions among pose landmarks. A decision threshold of 0.5 was applied to classify samples as stroke or non-stroke.
@@
Model performance was evaluated using Accuracy, Area Under the Receiver Operating Characteristic Curve (AUC), and the balanced F1-score. After data cleaning, the dataset consisted of 222 valid video samples and was split in a subject independent manner into training, validation, and test sets. Specifically, 29 subjects (116 videos) were used for training, while 4 subjects (24 videos) were allocated to each of the validation and test sets, ensuring no subject overlap across splits. All models were implemented in PyTorch and trained for 300 epochs using the Adam optimizer with the Binary Cross-Entropy loss function. All experiments were conducted on an NVIDIA GeForce RTX 3090 GPU.
V-B Experimental Result
@@
The results in the table III reveal a clear performance gap between unimodal and multimodal models, while also highlighting the different levels of contribution of each modality in the stroke screening task. Among the unimodal approaches, the pose-only model achieves the lowest performance across all metrics, with an Accuracy of 0.6667 and an F1-score of 0.6364. This indicates that upper-body pose information, when used in isolation, provides only limited motion-related cues and is insufficient to reliably distinguish between stroke and non-stroke conditions in the prehospital setting. The voice-only model shows a substantial improvement over the pose-only baseline, achieving both Accuracy and F1-score values of 0.8333. This suggests that speech signals contain more discriminative features associated with neurological impairments, particularly symptoms such as slurred or slowed speech. However, the Specificity of this model is limited to 0.8333, indicating the presence of a non-negligible number of false-positive predictions when relying solely on audio information. In contrast, the face-only model yields the highest performance among the unimodal approaches, with an Accuracy of 0.8750 and an AUC of 0.9306. Notably, this model achieves a Specificity of 1.0000, demonstrating its strong ability to correctly identify non-stroke cases. Nevertheless, its Sensitivity remains at 0.7500, implying that some stroke cases are still missed when relying exclusively on facial expressions and asymmetry cues.
@@
Table III: Performance comparison of different models. The best results are highlighted in bold.
@@
VI Ablation Study
Table IV presents a performance comparison of different fusion strategies to evaluate the impact of each feature aggregation method within the multimodal framework. Overall, learnable fusion strategies consistently outperform simple concatenation, indicating that explicitly modeling the relationships among modalities plays a crucial role in prehospital stroke screening.
@@ VI Ablation Study
Table IV: Performance Comparison of Different Fusion Strategies
@@
VII Dicussion
The experimental results demonstrate that integrating multimodal information provides clear advantages over unimodal approaches for prehospital stroke screening. Among individual modalities, facial and speech signals contribute more strongly than upper-body pose information. This finding aligns with the characteristics of the F.A.S.T. assessment, in which symptoms such as facial drooping and speech impairment are typically more salient than subtle upper-body motor abnormalities. The high Specificity achieved by the face-only model indicates that geometric facial features are effective in excluding non-stroke cases. However, its lower Sensitivity highlights the limitation of relying solely on facial expressions, as pronounced asymmetry may not always be present. In contrast, the voice-only model achieves a more balanced trade-off between Sensitivity and Specificity, suggesting that speech is a valuable diagnostic cue, though insufficient when used alone.
Multimodal fusion allows the model to exploit complementary information from facial appearance, speech, and motor behavior, resulting in a substantial improvement in discriminative performance. Notably, the fusion model with fine-tuned pretrained weights achieves the best overall performance across all metrics, indicating that adapting pretrained parameters to the target data facilitates stronger cross-modal interactions. Conversely, freezing pretrained weights improves training stability and AUC but does not fully optimize positive case detection.
@@ VII Dicussion
Overall, these results highlight the potential of multimodal learning combined with transfer learning in data-limited prehospital settings. Achieving a Sensitivity of 1.0000 with the fine-tuned fusion model is particularly important for early stroke screening, where minimizing missed cases is critical. At the same time, the high Specificity demonstrates reliable control of false alarms, supporting the feasibility of real-world deployment.
@@
Despite these promising results, several limitations related to data scale remain. The dataset is relatively small, consisting of 222 videos from a limited number of participants, which may restrict the model’s generalization to more diverse prehospital scenarios. In addition, the data were collected using simulated stroke symptoms rather than recordings from real patients. While this enables controlled acquisition and balanced labeling, it may not fully reflect the variability and complexity of real-world stroke presentations. The limited dataset size also increases the risk of overfitting, particularly for multimodal models with many parameters. Although pretrained models and controlled fine-tuning strategies improve stability, the reported performance should be interpreted cautiously and validated on larger, independent datasets.
VIII Conclusion

### v2

@@ 5.1 Implementation Details
The proposed system consists of multiple modality-specific branches designed to process heterogeneous input signals. Facial landmark sequences were modeled using a Transformer encoder initialized from a pretrained Audio Spectrogram Transformer (AST) backbone with a base configuration. Speech signals were processed using an AST model with a tiny configuration to extract discriminative acoustic representations from mel spectrograms. For upper-body pose data, an MLP-Mixer architecture was employed and trained from scratch to capture temporal and spatial interactions among pose landmarks.
All models were implemented in PyTorch and trained for 200 epochs using the Adam optimizer with the Binary Cross-Entropy loss function. A decision threshold of 0.5 was applied to classify samples as stroke or non-stroke. The experiments were conducted on an NVIDIA GeForce RTX 3090 GPU.
@@
After data cleaning, the dataset consisted of 222 valid video samples and was split in a subject-independent manner into training, validation, and test sets. Specifically, 29 subjects (116 videos) were used for training, while 4 subjects (24 videos) were allocated to each of the validation and test sets, ensuring no subject overlap across splits. Model performance was evaluated using Accuracy, Area Under the Receiver Operating Characteristic Curve (AUC), and the balanced F1-score. Table 3 summarizes the main model configurations and training hyperparameters used in our experiments.
[…]
Parameter Value
Face encoder AST (base)
Speech encoder AST (tiny)
Pose encoder MLP-Mixer
Embedding dimension 256
Optimizer Adam
Learning rate $1\times 10^{-4}$
Batch size 8
Training epochs 200
Loss function Binary Cross-Entropy
Decision threshold 0.5
Framework PyTorch
GPU NVIDIA RTX 3090
5.2 Computational Efficiency and Inference Cost
To address practical deployment considerations, we report the model size and inference efficiency of the proposed multimodal framework. Table 4 summarizes the parameter breakdown across modality-specific encoders and the fusion head. The model consists of three modality-specific branches and a lightweight fusion head. The face branch encodes temporal facial landmark sequences using a Transformer-based encoder (initialized from a pretrained AST encoder), which provides strong representation capacity but also contributes the majority of parameters. The speech branch converts raw audio into mel-spectrograms and applies an AST-tiny encoder to capture time-frequency patterns related to dysarthria; this branch is substantially smaller than the face encoder. The pose branch processes upper-body pose landmark sequences with an MLP-Mixer, offering a compact and efficient alternative to attention-heavy architectures while still modeling spatiotemporal motion cues. Finally, the fusion head aggregates the three embeddings via attention and produces the binary prediction with negligible parameter overhead. The overall model contains 92.490M parameters (352.82 MB when stored in 32-bit floating point), where the facial branch contributes the largest portion of parameters. In addition, we measure the end-to-end inference latency of the full multimodal pipeline at 7.19 ms per sample (batch size =4=4 during measurement). These results indicate that the proposed architecture can support real-time screening requirements, while also highlighting that future optimization efforts, such as model compression or quantization, can primarily focus on the facial encoder to further reduce the memory footprint.
[…]
Item Value
Face encoder params 85.791 M
Audio encoder params 5.809 M
Pose encoder params 0.643 M
Fusion + head params 0.247 M
Total params 92.490 M
Model size (params only) 352.82 MB
Inference time 7.19 ms / sample
5.3 Performance Comparison Across Model Variants
@@
The results in the table 5 reveal a clear performance gap between unimodal and multimodal models, while also highlighting the different levels of contribution of each modality in the stroke screening task. Among the unimodal approaches, the pose-only model achieves the lowest performance across all metrics, with an Accuracy of 0.6667 and an F1-score of 0.6364. This indicates that upper-body pose information, when used in isolation, provides only limited motion-related cues and is insufficient to reliably distinguish between stroke and non-stroke conditions in the prehospital setting. The voice-only model shows a substantial improvement over the pose-only baseline, achieving both Accuracy and F1-score values of 0.8333. This suggests that speech signals contain more discriminative features associated with neurological impairments, particularly symptoms such as slurred or slowed speech. However, the Specificity of this model is limited to 0.8333, indicating the presence of a non-negligible number of false-positive predictions when relying solely on audio information. In contrast, the face-only model yields the highest performance among the unimodal approaches, with an Accuracy of 0.8750 and an AUC of 0.9306. Notably, this model achieves a Specificity of 1.0000, demonstrating its strong ability to correctly identify non-stroke cases. Nevertheless, its Sensitivity remains at 0.7500, implying that some stroke cases are still missed when relying exclusively on facial expressions and asymmetry cues.
@@
To further evaluate the effectiveness of the proposed multimodal framework, we additionally compare our method with two commonly used multimodal baselines: Late Fusion and Feature Concatenation followed by a Multilayer Perceptron (Concat + MLP). The Late Fusion approach aggregates the predictions of the unimodal models by averaging their output probabilities. As shown in Table 5, this baseline achieves an Accuracy of 0.9167 and an F1-score of 0.9091, which is comparable to the best pairwise fusion configuration (Face + Voice). This result indicates that combining predictions from independently trained unimodal models can already capture complementary information across modalities to a certain extent. The Concat + MLP baseline, which directly concatenates the feature representations from the three modalities and feeds them into a multilayer perceptron for classification, achieves an Accuracy of 0.8750 and an F1-score of 0.8571. Although this approach allows the classifier to access multimodal information simultaneously, the simple concatenation strategy lacks an explicit mechanism to model complex cross-modal interactions. As a result, its performance remains comparable to some pairwise fusion configurations but does not fully exploit the complementary relationships among facial, speech, and pose features.
@@
[…]
@@
6 Ablation Study
6.1 Performance between difference fusion strategies
Table 6 presents a performance comparison of different fusion strategies to evaluate the impact of each feature aggregation method within the multimodal framework. In particular, we compare the proposed approach with several commonly used multimodal fusion baselines, including feature concatenation, simple summation, and learnable weighted summation. Overall, learnable fusion strategies consistently outperform simple concatenation, indicating that explicitly modeling the relationships among modalities plays a crucial role in prehospital stroke screening.
@@ 6.1 Performance between difference fusion strategies
[…]
@@
6.2 Performance under controlled modality corruptions
[…]
Corrupted Modality Performance
Face Pose Speech Acc AUROC F1 Sen Spec
0.9583 1.0000 0.9600 1.0000 0.9167
$\checkmark$ 0.9167 0.9861 0.9231 1.0000 0.8333
$\checkmark$ 0.9583 1.0000 0.9600 1.0000 0.9167
$\checkmark$ 0.8750 0.9722 0.8889 1.0000 0.7500
$\checkmark$ $\checkmark$ 0.9167 0.9861 0.9231 1.0000 0.8333
$\checkmark$ $\checkmark$ 0.8750 0.9722 0.8889 1.0000 0.7500
$\checkmark$ $\checkmark$ 0.8750 0.9583 0.8889 1.0000 0.7500
$\checkmark$ $\checkmark$ $\checkmark$ 0.8333 0.9514 0.8571 1.0000 0.6667
To assess robustness under challenging prehospital recording conditions, we perform a controlled test-time corruption study where each modality is corrupted while the remaining modalities are kept clean, and we additionally consider multi-modality corruptions. For Face and Pose, corruptions are applied at the raw video-frame level and landmarks are then re-extracted using the same MediaPipe pipeline. Specifically, we apply one of three visual perturbations to simulate common field conditions: (i) motion (random affine transformations to mimic handheld camera motion), (ii) outdoor (photometric changes via color jitter to emulate illumination/exposure variations), and (iii) blur (Gaussian blur to simulate defocus or motion blur). For Speech, we inject additive noise directly into the raw waveform at a fixed SNR (10dB) prior to mel-spectrogram computation, approximating moderate background noise. As reported in Table 7, the proposed multimodal model preserves perfect sensitivity across all corruption settings (Sen = 1.0000), which is critical for screening. Under single-modality corruptions, Pose corruption yields the smallest drop (Acc = 0.9583, F1 = 0.9600), Face corruption causes a moderate decrease (Acc = 0.9167, F1 = 0.9231), while Speech corruption shows the largest degradation among single-modality cases (Acc = 0.8750, F1 = 0.8889). When multiple modalities are corrupted simultaneously, performance decreases as expected; the most adverse setting (Face+Pose+Speech) results in Acc = 0.8333 and F1 = 0.8571, with the reduction mainly reflected in specificity (Spec = 0.6667), i.e., more false positives rather than missed positives, which is a preferable trade-off for a screening-oriented application.
6.3 Pose Branch Temporal-Segment Ablation
To further analyze the pose modality, we conduct a temporal-segment ablation on the upper-body pose landmark sequences. For each sample, we extract a fixed-length window of $T=200$ frames using three strategies: early (first $T$ frames), center (a centered window), and late (last $T$ frames). All other settings are kept unchanged.
Table 8 shows that the temporal window selection affects the pose-only performance. The center window yields the most balanced results, achieving Acc $66.67\%$ and F1 =0.6364=0.6364 with Sen =0.5833=0.5833 and Spec =0.7500=0.7500. The early window produces higher specificity (Spec =0.9167=0.9167) but substantially lower sensitivity (Sen =0.2500=0.2500), indicating more missed positive cases when only the initial segment is used. The late window achieves perfect specificity (Spec =1.0000=1.0000) but lower sensitivity (Sen =0.3333=0.3333), suggesting that while late-stage movements can be discriminative, relying only on the final segment is insufficient for consistent positive detection. Overall, these results indicate that coordination cues in the nose-touching task are not uniformly distributed over time, and the pose modality is sensitive to temporal cropping.
[…]
Window Acc AUROC F1 Sen Spec
Early 0.5833 0.7778 0.3750 0.2500 0.9167
Center 0.6667 0.8125 0.6364 0.5833 0.7500
Late 0.6667 0.8333 0.5000 0.3333 1.0000
Window TP FP FN TN
Early 3 1 9 11
Center 7 3 5 9
Late 4 0 8 12
7 Dicussion
@@
Overall, these results highlight the potential of multimodal learning combined with transfer learning in data-limited prehospital settings. Achieving a Sensitivity of 1.0000 with the fine-tuned fusion model is particularly important for early stroke screening, where minimizing missed cases is critical. At the same time, the high Specificity demonstrates reliable control of false alarms, suggesting that the proposed approach is promising as a screening-oriented decision-support prototype.
7.1 Simulated vs. Clinical Data
@@
Despite these promising results, several limitations related to data scale and clinical representativeness remain. The dataset is relatively small, consisting of 222 samples from a limited number of participants, which may restrict generalization to more diverse prehospital scenarios and increase the risk of overfitting, particularly for multimodal models with many parameters. To mitigate this limitation, we leverage transfer learning by initializing the visual and audio encoders from pretrained models, which helps improve representation learning under limited data conditions. In addition, data augmentation strategies were applied during training and analyzed in the ablation study to improve model robustness. In addition, the data were collected using simulated stroke symptoms rather than recordings from real patients; therefore, the current study should be regarded as a proof-of-concept rather than a clinically validated, deployable screening system. While simulated data enable controlled acquisition and balanced labeling, they may not fully reflect the variability and complexity of real-world stroke presentations (e.g., differences in symptom severity, mixed symptom profiles, comorbidities, and older demographics). Moreover, real prehospital recordings can involve uncontrolled conditions such as camera motion, occlusions, and background noise, which may introduce distribution shifts. Accordingly, validation on larger, independent datasets, ideally including external test sets and prospective evaluations on real patients, is required before deployment.
7.2 Demographic Bias and Age Generalizability
Our cohort is skewed toward younger participants (35/37 are aged 18–30), whereas stroke incidence is substantially higher in older populations. This age imbalance limits the current applicability of our findings, as age can change the baseline characteristics of cues used in F.A.S.T.-style screening. For example, facial geometry and habitual asymmetry may vary with age (e.g., sagging and reduced muscle tone), speech can be affected by age-related changes and comorbidities, and upper-body motion patterns may be influenced by reduced mobility or musculoskeletal conditions, potentially increasing false positives or masking subtle stroke-like deficits. Consequently, generalization to the target older population remains unproven in this study. In future work, we will expand data collection to include older participants and real patients, and perform age-stratified evaluation (and, if needed, calibration or domain adaptation) to ensure robust performance across age groups.
8 Conclusion